OpenStudy (anonymous):

A substance has a haft life of 1,523 years .What is fraction of the substance would be left after 10,000 years

OpenStudy (anonymous):

\[A _{1}=A _{0}e ^{kt}\]\[A=e ^{1/2(10,000/1523)}\] am i right with my formula?

OpenStudy (anonymous):

@jerwyn gayo No, it specifically says,half life so it goes like this: Substance left = \[\frac{1}{2}^{\frac{time}{halflife}}\] or, \[\frac{1}{2}^{\frac{1000}{1523}}\ \approx 0.01055\]

OpenStudy (anonymous):

@rlsadiz, you mean there is no (e)?